OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = 1/2x + 3.

OpenStudy (anonymous):

the slope of \(y=\frac{1}{2}x+3\) is \(\frac{1}{2}\) therefore the slope of a perpendicular line will be \(-2\)

OpenStudy (anonymous):

your job now is to use the "point - slope" formula \[y-y_1=m(x-x_1)\] with \[m=-2,x_1=-3,y_1=-1\]urmssmossymmh

OpenStudy (anonymous):

y = -2x - 7 ?

OpenStudy (anonymous):

i don't know i didn't do it. let me check

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

how did you do it